# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v5

@@ 5.4 Emergence During Training: How Divergent Strategies Develop
Two-Phase Representational Development. Conceptual organization emerges via two phases (Figure 8). First, rapid concept formation (1K-100K steps) establishes basic categorical structure, with AMI rising from near zero to approximately 0.45, achieving 80% of final alignment within just 10% of training. Second, architectural reorganization (100K-500K steps) systematically migrates semantic processing from deeper layers toward mid-network while AMI continues to gradually improve. This migration from layer 29 to layer 23 occurs without sacrificing categorical alignment, suggesting the model discovers more efficient internal representations.
@@ 7 Ethics statement
Our study relies exclusively on publicly available LLMs and digitized datasets from classic cognitive psychology experiments (Rosch, 1973a; 1975; McCloskey & Glucksberg, 1978). No new human subject data was collected, and all benchmark data we release have been properly attributed and curated to preserve research integrity.
@@ 8 Reproducibility statement
Appendix A Appendix
@@ Appendix B Limitations
B.1 Dataset Access Details
The aggregated and digitized human categorization datasets from Rosch (1973a; 1975); McCloskey & Glucksberg (1978) are made available in CSV format at: [URL deduced for anonymity; Data is attached as Supplementary Material].
B.2 LLM Details
  • BERT family: deberta-large, bert-large-uncased, roberta-large (Devlin et al., 2019; He et al., 2020; Zhuang et al., 2021).
  • QWEN family: qwen2-0.5b, qwen2.5-0.5b, qwen1.5-0.5b, qwen2.5-1.5b, qwen2-1.5b, qwen1.5-1.5b, qwen1.5-4b, qwen2.5-4b, qwen2-7b, qwen1.5-14b, qwen1.5-32b, qwen1.5-72b, qwen2.5-72b (Bai et al., 2023; Yang et al., 2024).
  • Llama family: llama-3.2-1b, llama-3.1-8b, llama-3-8b, llama-3-70b, llama-3.1-70b (Touvron et al., 2023a; b; Grattafiori et al., 2024).
  • Phi family: phi-1.5, phi-1, phi-2, phi-4 (Javaheripi et al., 2023; Abdin et al., 2024; Abouelenin et al., 2025).
  • Gemma family: gemma-2b, gemma-2-2b, gemma-7b, gemma-2-9b (Team et al., 2024; 2025).
  • Mistral family: mistral-7b-v0.3 (Karamcheti et al., 2021).
  • GPT family: gpt2, gpt2-medium (Radford et al., 2019).
  • OLMo family: olmo-7b (Groeneveld et al., 2024).
  • ViT family: Clip ViT-B/32, Clip ViT-B/16 (Radford et al., 2021).
  • Classic static embeddings: GloVe, Word2Vec (Pennington et al., 2014; Mikolov et al., 2013a; b).
B.3 Contextual Prompts and Pooling Strategies
Contextual embeddings of LLMs require feeding words into the model through a prompt. Because tokenizers often split a word into multiple tokens, and since some items in our datasets consist of two or more words, we face a design choice regarding how to aggregate token representations. In our methodology, we adopt average pooling over the actual tokens, ensuring that all subword pieces contribute equally. Figures 3 and 4 reveal that the average pooling strategy achieves consistent performance and demonstrates the tightest distribution, making it the most reliable choice for our research.
For prompts, we selected a neutral template, "This is a {word}. " (with a trailing space), designed to minimize any additional semantic bias on the target item. Figures 5 and 6 show this prompt to balance performance and consistency, making it ideal for baseline comparisons.
@@ B.3 Contextual Prompts and Pooling Strategies
B.4 Static vs. Contextual AMI Exploration
@@
B.5 Training Dynamics
@@
B.6 Datasets Polysemy
@@ B.7 Tokenizer Analysis
  • Qwen: Qwen1.5-0.5B (representative), Qwen1.5-1.8B, Qwen1.5-14B, Qwen2-0.5B, Qwen2-7B, Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-32B, Qwen1.5-32B, phi-4222phi-4 has a different tokenizer than the rest of Phi family. The results of its tokenizer match the tokenizer of the Qwen family.
B.8 Additional Clustering Metrics
@@
B.9 Detailed AMI Scores per Model and Dataset

### v6

@@ 5.4 Emergence During Training: How Divergent Strategies Develop
Two-Phase Representational Development. Conceptual organization emerges via two phases (Figure 15). First, rapid concept formation (1K-100K steps) establishes basic categorical structure, with AMI rising from near zero to approximately 0.45, achieving 80% of final alignment within just 10% of training. Second, architectural reorganization (100K-500K steps) systematically migrates semantic processing from deeper layers toward mid-network while AMI continues to gradually improve. This migration from layer 29 to layer 23 occurs without sacrificing categorical alignment, suggesting the model discovers more efficient internal representations. Moreover, this double-phase dynamics occurs when testing attention sparsity, effective rank, and $\mathcal{L}$ values. Meaning, all of which exhibit the same early rapid shift followed by a slower restructuring phase. This convergence across independent metrics indicates that the model is not merely improving categorical alignment, but reorganizing its internal representations toward increasingly efficient structure. See Appendix H.
@@ 7 Ethics statement
Our study relies exclusively on publicly available LLMs and digitized datasets from classic cognitive psychology experiments (rosch1973internal; rosch1975cognitive; mccloskey1978natural). No new human subject data was collected, and all benchmark data we release have been properly attributed and curated to preserve research integrity.
@@ 8 Reproducibility statement
Appendix A Cognitive Intuition
A.1 A Cognitive Intuition of the compression-meaning tradeoff
The compression-meaning tradeoff refers to the cognitive tension between representing concepts with maximal efficiency (i.e., minimal information) and preserving the semantic richness needed for flexible generalization, inference, and communication. For instance, upon hearing the sentence “There was a large brown Labrador barking loudly near the playground,” a person will often encode a simplified memory, such as “big scary dog near kids.” This is not to suggest that humans cannot recall the full sentence, but rather that we typically retain the most meaningful elements to enable efficient reasoning and generalization; without such abstraction and simplification, leveraging past experience for learning and prediction would be more difficult.
We acknowledge that our metrics, while capturing the information-compression trade-off and geometric efficiency in embedding space, do not directly measure all aspects of human-style conceptual abstraction or reasoning. Our aim is not to claim full equivalence with human cognition (nor do we think anyone can or should make such claims), but rather to provide a quantitative, interpretable proxy that highlights where LLMs and humans converge or diverge in how they compress and organize semantic information. We view these metrics as one lens among many, and we are careful in the paper to frame our findings as providing insights into human-like patterns rather than definitive evidence of human-style conceptual processing.
A.2 Cognitively-Inspired Inductive Biases
Future models could more closely align with human conceptual structure by incorporating cognitively motivated inductive biases and representational mechanisms. Hierarchical and compositional structure could enable models to capture nested relationships between categories, reflecting the way humans organize knowledge from superordinate to basic-level concepts (e.g., animal → mammal → dog → Labrador). Feature- and relation-based biases could help models focus on meaningful perceptual or functional attributes and relational patterns, rather than relying solely on statistical co-occurrence.
Additionally, theory- or causally grounded priors that draw on humans’ intuitive understanding of how objects interact or behave, could constrain learning in complex domains and support more flexible generalization. Incorporating a hybrid exemplar-rule approach, combining memory of specific examples with abstracted rules, would further approximate human category learning. Modular architectures, in which specialized sub-networks handle different aspects of conceptual representation, could enhance generalization and reduce interference between unrelated features. Finally, meta-learned priors distilled from symbolic or program-based representations offer a way to embed structured, human-like concept hypotheses directly into neural models, allowing them to generalize more like humans across novel situations.
Together, these inductive biases offer a path for models that not only compress information efficiently but also organize knowledge in a manner that mirrors human conceptual richness, capturing graded typicality, family resemblance, and hierarchical relationships. While integrating these biases may trade some compression for fidelity, they provide an exciting opportunity to reduce meaning distortion and bridge the gap between statistical efficiency and human-like conceptual understanding.
@@ Appendix B Limitations
Appendix C Dataset Access Details
The aggregated and digitized human categorization datasets from rosch1973internal; rosch1975cognitive; mccloskey1978natural are made available in CSV format at: [URL deduced for anonymity; Data is attached as Supplementary Material].
Appendix D LLM Details
  • BERT family: deberta-large, bert-large-uncased, roberta-large (devlin-etal-2019-bert; he2020deberta; zhuang-etal-2021-robustly).
  • QWEN family: qwen2-0.5b, qwen2.5-0.5b, qwen1.5-0.5b, qwen2.5-1.5b, qwen2-1.5b, qwen1.5-1.5b, qwen1.5-4b, qwen2.5-4b, qwen2-7b, qwen1.5-14b, qwen1.5-32b, qwen1.5-72b, qwen2.5-72b (bai2023qwen; yang2024qwen2).
  • Llama family: llama-3.2-1b, llama-3.1-8b, llama-3-8b, llama-3-70b, llama-3.1-70b (touvron2023llama; touvron2023llama2; grattafiori2024llama).
  • Phi family: phi-1.5, phi-1, phi-2, phi-4 (javaheripi2023phi; abdin2024phi; abouelenin2025phi).
  • Gemma family: gemma-2b, gemma-2-2b, gemma-7b, gemma-2-9b (team2024gemma; team2025gemma).
  • Mistral family: mistral-7b-v0.3 (karamcheti2021mistral).
  • GPT family: gpt2, gpt2-medium (radford2019language).
  • DeepSeek family: DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B, DeepSeek-R1-Distill-Qwen-32B, DeepSeek-R1-Distill-Llama-8B (guo2025deepseek).
  • OLMo family: Olmo-7b (olmo2024).
  • ViT family: Clip ViT-B/32, Clip ViT-B/16 (radford2021learning).
  • Classic static embeddings: GloVe, Word2Vec (pennington-etal-2014-glove; mikolov-etal-2013-efficient; mikolov-etal-2013-distributed).
Appendix E Contextual Prompts and Pooling Strategies
Contextual embeddings of LLMs require feeding words into the model through a prompt. Because tokenizers often split a word into multiple tokens, and since some items in our datasets consist of two or more words, we face a design choice regarding how to aggregate token representations. In our methodology, we adopt average pooling over the actual tokens, ensuring that all subword pieces contribute equally. Figures 4 and 5 reveal that the average pooling strategy achieves consistent performance and demonstrates the tightest distribution, making it the most reliable choice for our research.
For prompts, we selected a neutral template, "This is a {word}. " (with a trailing space), designed to minimize any additional semantic bias on the target item. Figures 6 and 7 show this prompt to balance performance and consistency, making it ideal for baseline comparisons.
@@ Appendix E Contextual Prompts and Pooling Strategies
Appendix F Static vs. Contextual AMI Exploration
@@
We also tested the robustness of our results to different clustering seeds (Figure 9). We found AMI to be highly stable across seeds, with negligible variation in the peak values and layer-wise profiles, indicating that our conclusions are not sensitive to the choice of clustering initialization.
Figure 9: AMI Scores are Robust Across Clustering Seeds. Peak AMI between model representations and human categories remains highly stable across multiple random clustering initializations. Peaks in each plot are consistent, indicating that the observed alignment patterns are not sensitive to the choice of clustering seed.
Lastly, we plot the peak AMI against the number of FLOPS per token and find no systematic correlation, suggesting that computational cost alone does not predict human-aligned conceptual representations (Figure 10.)
Figure 10: Peak AMI versus computational cost. We plot each model’s peak AMI against its FLOPS per token. There is no systematic correlation, indicating that higher computational cost does not necessarily lead to better alignment with human conceptual structure.
Appendix G Multilingual Analysis
To further explore conceptual understanding across languages, we translated our dataset into Spanish, German, Italian, and Russian using Google Translate’s API and repeated our analyses with the same LLMs and methods. In RQ1, a clear scale effect emerges for all non-English languages, while English shows no such trend (Figures 11, 12). We interpret this as a consequence of limited non-English training data: larger models are more likely to have been exposed to sufficient multilingual data, improving their conceptual alignment. In RQ2, all models struggle to preserve the internal geometry of human concepts across languages (Figure 13). RQ3 shows that non-English languages exhibit greater compression (Figure 14), consistent with our explanation for RQ1: smaller exposure to non-English data leads to more compressed representations, reducing flexibility and interpretability.
Figure 11: Scale effect emerges for non-English languages. We compute peak AMI as a function of model size across different languages: English (red), Spanish (blue), Italian (purple), German (yellow), and Russian (pink). While English shows no systematic scaling effect, the other languages exhibit a clear positive relationship between model size and AMI. We interpret this as a consequence of limited non-English training data: larger models are more likely to have been exposed to sufficient multilingual data, improving their conceptual alignment. Additional analyses across other RQs using the multilingual data support this hypothesis.
Figure 12: Scale effect emerges for non-English languages. We compute peak AMI as a function of model size across different languages: English (red), Spanish (blue), Italian (purple), German (yellow), and Russian (pink). While English shows no systematic scaling effect, the other languages exhibit a clear positive relationship between model size and AMI. We interpret this as a consequence of limited non-English training data: larger models are more likely to have been exposed to sufficient multilingual data, improving their conceptual alignment. Additional analyses across other RQs using the multilingual data support this hypothesis.
Figure 13: Preservation of internal conceptual geometry across languages. Models struggle to maintain the internal structure of human concepts in all tested languages, with similar patterns observed across English, Spanish, German, Italian, and Russian.
(a) $\mathcal{L}$ as a function of K for rosch1973internal.
(b) $\mathcal{L}$ as a function of K for rosch1975cognitive.
(c) $\mathcal{L}$ as a function of K for mccloskey1978natural.
Figure 14: Compression of non-English representations. All non-English languages (Spanish, German, Italian, Russian) exhibit higher compression than English, with smaller models showing the greatest compression. This supports the interpretation that limited non-English training data leads to less flexible and less interpretable representations.
Appendix H Training Dynamics
@@
Moreover, this double-phase dynamics occurs when testing attention sparsity, effective rank, and $\mathcal{L}$ values (Figures 18, 19). Meaning, all of which exhibit the same early rapid shift followed by a slower restructuring phase. This convergence across independent metrics indicates that the model is not merely improving categorical alignment, but reorganizing its internal representations toward increasingly efficient structure.
@@
(a) Effective rank across network layers.
(b) Gini coefficient across network layers.
(c) Hoyer sparsity across network layers.
Figure 18: Network-level measures of representational structure across training. Each panel shows a different metric: (a) effective rank, (b) Gini coefficient, and (c) Hoyer sparsity computed across network layers. All three measures reveal the same two-phase developmental pattern observed in Figure 15: an early rapid change followed by slower restructuring, indicating coordinated reorganization of internal representations.
Figure 19: $\mathcal{L}$ across training. Across all three human benchmarks, Olmo training checkpoints reveal the same two-phase developmental pattern observed in Figure 15: an early rapid change followed by slower restructuring, indicating coordinated reorganization of internal representations.
Appendix I Datasets Polysemy
@@
  • Qwen: Qwen1.5-0.5B (representative), Qwen1.5-1.8B, Qwen1.5-14B, Qwen2-0.5B, Qwen2-7B, Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-32B, Qwen1.5-32B, phi-4333phi-4 has a different tokenizer than the rest of Phi family. The results of its tokenizer match the tokenizer of the Qwen family.
Appendix K Additional Clustering Metrics
@@
Appendix L Mini-Controlled Experiment (Matched Training Data)
To evaluate the extent to which dataset differences might account for the architectural patterns we report, we conducted matched-family analyses involving the only model families that can be aligned in both training data: GPT, Pythia, Cerebras, and T5. While these comparisons cannot rule out all confounds, they substantially reduce the influence of dataset variation. Across all matched settings, encoder models continue to outperform decoder models, yielding higher AMI and lower $\mathcal{L}$. This indicates that the architectural effects observed throughout the paper cannot be explained by differences in training data alone.
(a) Peak AMI as a function of model size.
(b) Layer-wise Spearman correlation structure.
Figure 24: Evidence that encoder-decoder differences are not driven by dataset artifacts. We compare the only model families that can be matched in training data GPT, Pythia, Cerebras, and T5. While these comparisons cannot eliminate all possible confounds, they show that the architectural patterns we report are robust: encoder models consistently achieve higher AMI and lower $L$, suggesting that the observed differences cannot be explained by dataset variation alone.
Appendix M Detailed AMI Scores per Model and Dataset